log5x=log4+log(x+1)
Answer by nerdybill(2694)

You will need to apply "log rules". For a refresher see:
http://www.purplemath.com/modules/logrules.htm
.
log5x=log4+log(x+1)
log5x=log(4(x+1))
log5x=log(4x+4)
5x=4x+4
x=4